Search for an average annual salary of an engineer with the specialization that you are studying...

  1. Search for an average annual salary of an engineer with the specialization that you are studying (Mechanical Engineering) online (internet).
  2. Make a list of the expenses and their quantities that the engineer will have in any given year in the practice of his or her profession. Not in his or her personal life.
  3. Compute the taxes and the after tax cash flow for that year.
  4. Also show in where did you find the average annual salary and how you determine the expenses.
Answer #1

Average Annual Salary of a Mechanical Engineer in US: $85,880

Tax paid @ 24% when under bracket $84,201 to $160,725: $20,611

Total Left over post tax: $65,269

Expenses for the Year:

1. Avg. Cost of Engineering (4 years): $169,660

2. EMI for Education/Student Loan of $169,000 @ 5% for 6 years: $2,657 * 12 = $31,884

3. Any external certification: $5,000

4. Plan for Higher Studies Preparation: $1,000

Total left over post professional experience: 65269 - 31884 - 5000 - 1000 = $27,385

Per month left over: $27,385 / 12 = $2,282

$2,282 is Good for personal expense and savings.

Hope this helps. In case you find it helpful then give a thumbs up.